OpenStudy (legomyego180):

Whats the easiest way to solve for x here? 4sqrt(x)=(1/16)x^2

OpenStudy (legomyego180):

\[4\sqrt{x}=\frac{ 1 }{ 16 }x^2\]

OpenStudy (legomyego180):

im doing this for integration to find the area between the curves, and it seems like completing the square for this is a really long process when added on top of the integration. I bet there's a trick to know the bounds faster

OpenStudy (shawn):

graphing?

OpenStudy (shawn):

wolfram alpha. Can't get faster than that lol

OpenStudy (peachpi):

maybe factoring if you have to do it by hand \[4\sqrt{x}-\frac{1}{16}x^2=0\] \[\frac{ 1 }{ 16 }\sqrt{x}(64-x^\frac{3}{2})=0\] \[\sqrt{x}=0 \rightarrow x=0\] \[(64-x^\frac{3}{2})=0\rightarrow x=16\]

OpenStudy (legomyego180):

I like that @peachpi I didnt think of that thank you. Wolfram Alpha is great but unfortunately I cant use it on the test :P